OpenStudy (anonymous):

y+2x=5;x=-1,0,3 what is the value of y and how do you do it

OpenStudy (anonymous):

okay so lets start off by saying that you need to isolate your varaible of y. to do this you would subtract the +2x and move it to the other side. From here it should be more straight forward because you would just plug in your variables to find your solutions/values for y y = -2(-1)+5 y= -2(0)+5 y= -2(3)+5

OpenStudy (anonymous):

can you find your values of y now ? (: